OpenStudy (anonymous):

The formula P = ns gives the formula for the perimeter of a regular polygon with n sides and side length s. What is the length of each side of a regular octagon with a perimeter of 54.4 cm? A regular octagon has 8 equal sides.

sammixboo (sammixboo):

Well plug in the info :) \(\rm \color{#20bd23}{P~=~ns}\) p = perimeter = 54.4 n = number of sides = 8 s = side length = unknown \(\rm \color{#20bd23}{54.4~=~8s}\) To solve this, all you need to do is divide 8 on both sides of the equation. \(\rm \color{#20bd23}{\dfrac{54.4}{8}~=\dfrac{8s}{8}}\) \(\rm \color{#20bd23}{\dfrac{54.4}{8}~=~s}\)

sammixboo (sammixboo):

So what is \(\color{#20bd23}{54.4\div 8}\)?

OpenStudy (anonymous):

6.8